OpenStudy (anonymous):

A boat is traveling up river from New Orleans to Baton Rouge (90 miles) returning from Mardi Gras. The boat can travel at 22 mph in still water, and the Mississippi River has a flow rate of 9 mph. If the boat needs to arrive in Baton Rouge at 4:00 PM, what time does it need to leave New Orleans?

OpenStudy (kropot72):

Would you like to go thru the solution with me?

OpenStudy (anonymous):

Yes Please. I am mainly having trouble setting up an equation.

OpenStudy (kropot72):

I sure we can set up an equation. The actual speed of the boat is S(actual) = (still water speed) - (flow rate of river) Can you put the numbers in?

OpenStudy (anonymous):

Yes. I got 22= x - 9 ?

OpenStudy (kropot72):

You don't need a variable at this stage. S(actual speed of boat) = 22 - 9 = ? mph

OpenStudy (anonymous):

Okay, 13 mph is the actual speed of the boat. Alright that helps a lot just having those variables figured out. Thank You!

OpenStudy (anonymous):

so would the boat have to leave at 8:55 am?

OpenStudy (kropot72):

So the boat has to travel 90 miles at 13 mph. Therefore it will take 90/13 hours = 6 hours and 55 minutes (approximately). The boat will need to leave at 9:05 am.

OpenStudy (anonymous):

awesome thank you!

OpenStudy (kropot72):

You're welcome :)